# Simplifying generic smooth maps to the $2$–sphere and to the plane

### v1

@@ 7. Elimination of indefinite folds of certain indices
Let us cut open $W_{f}$ along $N(\alpha_{s})$, $s=1,2,\ldots,r$ so that we obtain the bottom figure of Fig. 19. Let $\beta$ be a properly embedded arc in $Y$ as in the figure. First, note that for each $s$, $M_{s}=q_{f}^{-1}(\alpha_{s})$ is a closed connected $(n-1)$–dimensional manifold and that $q_{f}|_{M_{s}}:M_{s}\to\alpha_{s}\cong[-1,1]$ is a Morse function with exactly one minima and one maxima and without critical points of index $1$ or $n-2$. Hence $M_{s}$ is simply connected. Let $h:Y\to\mathbb{R}$ be a smooth function which corresponds to the negative of the hight function as depicted in the bottom figure of Fig. 19. Then we may assume that $h\circ q_{f}:q_{f}^{-1}(Y)\to\mathbb{R}$ is a Morse function and it has exactly $r-1$ critical points of index $1$, each of which corresponds to a $1$–handle connecting $M_{s_{1}}$ and $M_{s_{2}}$, $s_{1}\neq s_{2}$. Hence, $q_{f}^{-1}(\beta)$ is diffeomorphic to the connected sum $M_{1}\sharp M_{2}\sharp\cdots\sharp M_{r}$, which is simply connected.
@@
This implies that the inclusion induces an isomorphism $H_{i}(F)\to H_{i}(X)$ for $i<(n-3)/2$. Thus, the inclusions $X\to X_{\pm}\simeq F$ also induce isomorphisms in homology. Hence, by (7.1) we have the short exact sequence
@@
This implies that the inclusion induces an isomorphism $H_{i}(F)\to H_{i}(M)$. Since $B$ and $F$ are simply connected by Lemma 7.1 and $M$ is $k$–connected, we see that $B$ and $F$ are also $k$–connected. This completes the proof. ∎
@@
The author [15] showed that a closed orientable $3$–dimensional manifold $M$ admits an image simple generic map $f:M\to N$ without cusp points for a surface $N$ if and only if $M$ is a graph manifold. Note that a closed orientable $3$–dimensional manifold is a graph manifold if it is the union of some $S^{1}$–bundles over compact surfaces attached along the torus boundaries. Note that the class of graph manifolds is relatively small in the set of closed orientable $3$–dimensional manifolds: for example, hyperbolic $3$–manifolds never admit such an image simple generic map.
@@ Acknowledgment
The author would like to thank Professor İnanç Baykur for stimulating discussions which motivated the theme of this paper. He would also like to thank Dominik Wrazidlo and Rustam Sadykov for some discussions. The author would also like to thank Professor Vincent Blanlœil and the Advanced Mathematical Research Institute of the University of Strasbourg for the hospitality during the preparation of the manuscript. This work has been supported in part by JSPS KAKENHI Grant Numbers JP22K18267, JP23H05437. This work was also supported by the Research Institute for Mathematical Sciences, an International Joint Usage/Research Center located in Kyoto University.

### v2

@@ 7. Elimination of indefinite folds of certain indices
Let us cut open $W_{f}$ along $N(\alpha_{s})$, $s=1,2,\ldots,r$, so that we obtain the bottom figure of Fig. 20. Let $\beta$ be a properly embedded arc in $Y$ as in the figure. Let us consider the structure of $q_{f}^{-1}(\beta)$. First, note that for each $s$, $M_{s}=q_{f}^{-1}(\alpha_{s})$ is a closed connected $(n-1)$–dimensional manifold and that $q_{f}|_{M_{s}}:M_{s}\to\alpha_{s}\cong[-1,1]$ is a Morse function with exactly one minimum and one maximum and without critical points of index $1$ or $n-2$. Hence $M_{s}$ is simply connected. Let $h:Y\to\mathbb{R}$ be a smooth function which corresponds to the negative of the height function as depicted in the bottom figure of Fig. 20. Then we may assume that $h\circ q_{f}:q_{f}^{-1}(Y)\to\mathbb{R}$ is a Morse function. Note that its critical points correspond to those of $h\circ q_{f}$ restricted to the singular point set $S(f)\cap q_{f}^{-1}(Y)$, and we may assume that the cusp points are not critical points. Since $f$ does not have folds of absolute index $1$, we see that the critical points on indefinite folds have indices different from $10,1,n-$ 1 or $n$. Furthermore, the function has a unique critical point of index $n$ together with exactly $r-1$ critical points of index $1$, each of which corresponds to a $1$–handle connecting $M_{s_{1}}$ and $M_{s_{2}}$, $s_{1}\neq s_{2}$. Hence, $q_{f}^{-1}(\beta)$ is diffeomorphic to the connected sum $M_{1}\sharp M_{2}\sharp\cdots\sharp M_{r}$, which is simply connected.
@@
Let us now consider elimination of folds with higher absolute indices. For this, we need the following.
@@
This implies that the inclusion induces an isomorphism $H_{i}(F)\to H_{i}(X)$ for $i<(n-3)/2$. Thus, the inclusions $X\to X_{\pm}\simeq F$ also induce isomorphisms in homology for dimensions $<(n-3)/2$. Hence, by (7.1) we have the short exact sequence
@@
for $i<(n-3)/2$. This implies that the inclusion induces an isomorphism $H_{i}(F)\to H_{i}(M)$ for $i<(n-3)/2$. Since $B$ and $F$ are simply connected by Lemma 7.1 and $M$ is $k$–connected with $k<(n-3)/2$, we see that $B$ and $F$ are also $k$–connected by the Hurewicz theorem. This completes the proof. ∎
@@
Proof.
By Theorem 7.2, we can eliminate folds of absolute index $1$ with the expense of creating folds of absolute indices $2$ and $3$. As $F$ is $k$–connected, we can use the same argument to successively eliminate folds of absolute indices $2,3,\ldots,k$ with the expense of creating folds of absolute indices $k+1$ and $k+2\leq(n-1)/2$. This completes the proof. ∎
@@
In the proofs of Theorems 7.2 and 7.6, we have constructed generic smooth maps of $M$ into $\mathbb{R}^{2}$ without cusps which are compatible with the open book structure constructed in the proof of Theorem 6.3.
Remark 7.8.
We do not know if the condition $21\leq k\leq(n-5)/$ 2 can be weakened or not. For example, we do not know if a $1$–connected $5$–dimensional manifold always admits a generic map into $\mathbb{R}^{2}$ without cusps and with only folds of absolute indices 00 or $2$.
Remark 7.9.
In Theorem 7.6, the condition that $M$ should be $k$–connected is not a necessary condition for the existence of a stable map $f:M\to\mathbb{R}^{2}$ without cusp points such that $f$ has no fold points of absolute index $i$ with $1\leq i\leq k$. As the proof of Lemma 7.3 shows, even if such an $f$ exists, $M$ may have nontrivial free fundamental group. For example, there exist a plenty of explicit examples of manifolds with nontrivial free fundamental groups which admit generic maps into $\mathbb{R}^{2}$ with only folds of absolute index 00 (see [20], for example).
Remark 7.10.
We do not know if in Theorem 7.6 the condition that $n\geq 7$ is odd and $21\leq k\leq(n-5)/$ 2 is essential or not.
8. Singular point set
In this section, we consider the topological position of the singular point set of a generic map, instead of its image. Such study was done in the author’s previous paper [21]; however, in some cases, the result was not correctly stated. In this section, we give a correction to the result.
More precisely, in [21], the author has studied the position of the singular point set $S(f)$ of a generic map $f:M\to N$ of a closed $n$–dimensional manifold, $n\geq 3$, into a surface $N$. However, when $n$ is even, the main result is not correctly stated and the proof needs also to be modified appropriately, as follows.
Let $f:M\to N$ be a generic map of a closed $n$–dimensional manifold into an oriented surface $N$, where we assume that $n\geq 2$ is even. In this case using the map $f|_{S(f)}$, we can naturally orient $S(f)$ as follows. Let $C$ be a component of $S(f)\setminus\{\mbox{cusps}\}$ with absolute index $i$. Recall that $f|_{C}:C\to N$ is a normally oriented immersion. For a point $p\in C$, we orient $T_{p}C$ so that the ordered basis of $T_{f(p)}N$ consisting of a normal orientation and the image by $df_{p}$ of a nonzero vector in $T_{p}C$ consistent with the orientation gives the orientation (or opposite orientation) of $T_{f(p)}N$ when $i$ is even (resp. odd). Then, with the help of Fig. 3, we see that this gives a unique orientation for each component of $S(f)$. In the rest of this section, we always consider this orientation for $S(f)$.
Note that such an orientation cannot be defined canonically when $n$ is odd.
Then, the main result of [21] for the case of $n$ even, should be stated as follows.
Theorem 8.1.
Suppose that $M$ is a closed $n$–dimensional manifold with $n\geq 2$ even and that $N$ is an oriented surface. Let $g:M\to N$ be a generic map. Then, for any nonempty closed oriented $1$–dimensional submanifold $L$ of $M$ which is $\mathbb{Z}$–homologous to $S(g)$, there exists a generic map $f:M\to N$ homotopic to $g$ such that $S(f)=L$.
Proof.
The proof is almost the same as that given in [21]; however, we need to be careful with the orientations of $S(g)$ and $L$.
Recall the (oriented) band operation introduced in [21, Definition 3.6]. As the proof of [21, Lemma 3.9] shows, $L$ is oriented isotopic to a $1$–dimensional oriented submanifold of $M$ obtained from $S(g)$ by a finite iteration of oriented band operations.
In the proof of [21, Theorem 2.2] for $n$ even, just after [21, Figure 12], we chose a joining curve $\lambda$ connecting $p_{1}$ or $p_{2}$ and another cusp in $S(g)$. In fact, both of $p_{1}$ and $p_{2}$ can be used, and there is no difference. This is because we can easily observe that the oriented band operations associated with $\lambda$ for $p_{1}$ and for $p_{2}$ give the same result, considering the orientations. Then, in order to adjust the information mentioned in [21, Remark 3.11 (2)], it has been claimed that choosing $p_{1}$ or $p_{2}$ appropriately, one can adjust it; however, as they do not have any difference, this does not work, unfortunately. Therefore, we need the information on the orientations. If the band is already oriented, then the relevant data is automatically adjusted.
Then, the rest of the proof works well without changing anything. This completes the proof. ∎
@@ Acknowledgment
The author would like to thank Professor İnanç Baykur for stimulating discussions which motivated the theme of this paper. He would also like to thank Dr. Rustam Sadykov for some important comments about certain moves and for essential discussions. He would also like to thank Dr. Dominik Wrazidlo for some discussions. The author would also like to thank Professor Vincent Blanlœil and the Advanced Mathematical Research Institute of the University of Strasbourg, and also to the Department of Mathematics, Kansas State University, for the hospitality during the preparation of the manuscript. This work has been supported in part by JSPS KAKENHI Grant Numbers JP22K18267, JP23H05437. This work was also supported by the Research Institute for Mathematical Sciences, an International Joint Usage/Research Center located in Kyoto University.